dudeIDK:

What is the rule for the function shown in the table? A. y = one-thirdx + 1 B. y = 3x + 1 C. y = one-thirdx D. y = one-halfx + 1

jacksonsl:

um

jacksonsl:

its not d are a

jacksonsl:

so choose one in between

jacksonsl:

C or B

dudeIDK:

thx

4RCH4N63L:

Today's video is sponsored by Raid Shadow Legends, one of the biggest mobile role-playing games of 2019

dudeIDK:

it 2021 but i do agree the game is good

ScamTheMan:

Can you post the table.

dudeIDK:

huh

dudeIDK:

oh ye

ScamTheMan:

The question is asking for the function represented in the table. You didn't post a table

dudeIDK:

X -1 0 1 2 y -2 1 3 7

YSNJayy:

Today's video is sponsored by Raid Shadow Legends, one of the biggest mobile role-playing games of 2019 and it's totally free! Currently almost 10 million users have joined Raid over the last six months, and it's one of the most impressive games in its class with detailed models, environments and smooth 60 frames per second animations! All the champions in the game can be customized with unique gear that changes your strategic buffs and abilities! The dungeon bosses have some ridiculous skills of their own and figuring out the perfect party and strategy to overtake them's a lot of fun! Current

4RCH4N63L:

wth

dudeIDK:

already have the game people

4RCH4N63L:

bro thats my line

4RCH4N63L:

Today's video is sponsored by Raid Shadow Legends, one of the biggest mobile role-playing games of 2019 and it's totally free! Currently almost 10 million users have joined Raid over the last six months, and it's one of the most impressive games in its class with detailed models, environments and smooth 60 frames per second animations! All the champions in the game can be customized with unique gear that changes your strategic buffs and abilities! The dungeon bosses have some ridiculous skills of their own and figuring out the perfect party and strategy to overtake them's a lot of fun! Currently with over 300,000 reviews, Raid has almost a perfect score on the Play Store! The community is growing fast and the highly anticipated new faction wars feature is now live, you might even find my squad out there in the arena! It's easier to start now than ever with rates program for new players you get a new daily login reward for the first 90 days that you play in the game! So what are you waiting for? Go to the video description, click on the special links and you'll get 50,000 silver and a free epic champion as part of the new player program to start your journey! Good luck and I'll see you there!

YSNJayy:

WOOOOOOOOOOWWWww

4RCH4N63L:

i am gonna get banned

YSNJayy:

you gon take my light

YSNJayy:

mods arent on moron

ScamTheMan:

Well, to figure out answer to the question, you want to insert your numbers into your formulas and see which one works out. For answer choice A we just substitute like so: \[y=1/3x +1\] \[1=1/3(0)+1\] As you can see, this is clearly true, however we have only used one value from the table. Using another, such as (2,7): \[7=1/3(2) + 1\] We can see that this is not the correct answer choice. Simply said, to figure out the answer, you just substitute.

ScamTheMan:

Now as we try out the equations to see if they're correct, we can see that option B.), makes the most sense. Time to try out the points in the table to see if this is the answer: \[y=3x+1\] \[1=3(0)+1\] \[7=3(2)+1\] As you can see, these are both true, therefore we can assume this is the answer. Plus the other answer choices represent x with a fraction, meaning the y value wouldn't exceed the x by such a large margin.

dudeIDK:

thx

ScamTheMan:

No problem, you understand how to do it?

dudeIDK:

ye

ScamTheMan:

Lovely, make sure to "Best Response" as well And if you need further help, make sure to ask!

xxDeppressionxx:

\(\style{fontFamily:Comic Sans;text-shadow: 1px 1px 2px black, 0 0 10px #00f7e6, 0 0 5px darkblue;color:black;font-size:100px;font-weight:bold;padding:50px;width:500px;height:75%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\sf\color {} {{Yeah ~you~can~ask~anyone~here}}}\)

jhonyy9:

good job @ScamTheMan

ScamTheMan:

Thank you <3

jhonyy9:

yw